### v1

# Role of sub-barrier dynamics in coherent and incoherent radiation during tunneling ionization

Abstract
Radiation during strong-field tunneling ionization is investigated. The spontaneous as well as the coherent component of the radiation are calculated and the role of the under-the-barrier dynamics is analyzed. The competition of the contribution of different mechanisms in different spectral regions is examined. Transition-like radiation (Brunel radiation) is dominated at low-frequencies, Thomson scattering – at the laser frequencies, and radiative recombination via the three-step process – at high-order harmonics. The under-the-barrier dynamics is shown to enhance the Brunel radiation for near-zero-frequencies in asymmetric laser pulses for spontaneous as well as for the coherent radiation.
I Introduction
It is well known that tunneling ionization is accompanied by high-order harmonic generation (HHG) via the three-step model [1], when the recombination of the ionized electron wave packet to the bound state at the recollision after the excursion in the laser field and gaining energy results in the emission of a high energy photon [2, 3, 4]. At the recollision, radiation may emerge also due to electron scattering, rather than recombination, yielding free-free HHG [5]. In the macroscopic picture, the ionization of a gas target results in the creation of an electron current after the interaction with a laser pulse, which can be a source of low frequency THz radiation, with the emitted frequency determined by the laser pulse envelope [6, 7, 8, 9]. The same mechanism can work also in the case of a single atom producing THz radiation in asymmetric laser pulses, e.g. in a two-color laser field [10, 11].
In contrast to the coherent radiation discussed above, emission of spontaneous low-order harmonic radiation during the bound-continuum transition with the origin similar to Brunel radiation [12, 13] has been recently shown in Ref. [14], with the typical frequency determined by the Keldysh time [15]. Finally, harmonic radiation at tunneling ionization arises also from Thomson scattering of the laser wave off the ionized free electron wave packet, with higher harmonics than the fundamental one suppressed in the nonrelativistic regime.
Coherent radiation should be distinguished from the incoherent one of the same origin. For instance, the spontaneous recombination of a high-energy electron to an atomic bound state results in a photon emission of a random phase, which is an incoherent process [16]. However, when the recombining electron is released from the bound state and recombines into the same bound state driven by a laser field, the HHG is coherent with the phase determined by the coherent superposition of the ground state wave function with the ionization continuum wave packet. Similarly, the coherent part of radiation due to Compton scattering of a strong laser wave from an ionized electron wave packet [17, 18] is phase-matched in forward direction, while the incoherent one has a broad angular distribution [19, 20, 21, 22, 23]. For the incoherent process, in contrast to the coherent one, the final state of the electron does not exist initially. The coherent radiation intensity scales as $\sim{\cal N}^{2}$ with the number of emitters $\cal N$}, due to the phase matching of the radiation, while for the incoherent radiation interference is suppressed and the total intensity scaling is $\sim{\cal N}$ [24]. The coherent radiation can be calculated using the quantum mechanical expectation value of the current density within the classical radiation theory [19].
In this paper we investigate the role of the under-the-barrier dynamics for the spontaneous and coherent Brunel radiation during tunneling ionization. We analyze how the different mechanisms of the radiation during the strong field ionization are scaled in different spectral regions of emission and the feasibility of their observation in a coherent or incoherent form.
The harmonics emitted from a gas target is the result of a multi-atom contribution, and therefore, are dominated by the coherent radiation. For the observation of the incoherent radiation either a single atom experiment should be carried out (when each time a single atom existed in the interaction region), or the emission should be measured at a stray direction far from the phase-matching one [22].
The structure of the paper is the following. In Sec. II the coherent radiation during tunneling ionization is calculated using the strong field approximation (SFA). Spontaneous emission at the tunneling ionization is discussed in Sec. III. We give estimations of radiation intensities in Sec. IV. Our conclusion is given in Sec. V.
II Coherent radiation during tunneling ionization
Figure 1: Coherent radiation spectra during tunneling ionization with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=6$: (a) in the asymmetric laser pulse, (b) in the symmetric laser pulse. Comparison of the Brunel radiation via $j_{11}$ with the three-step HHG ($j_{01}+j_{10}$) and with the Babushkin model ($j_{B}$). The total radiation with the full current $j=j_{01}+j_{10}+j_{11}$ is given for comparison.
We calculate coherent radiation emitted during tunneling ionization of a laser-driven hydrogenlike atom for the nonrelativistic regime. The dynamics is described by the Schrödinger equation $i\partial_{t}=H$ with Hamiltonian
$H=\mathbf{p}^{2}/2+V(r)+\mathbf{r}\cdot\mathbf{E}(t)+\mathbf{r}\cdot\hat{\mathbf{E}}(t),$ (1)
where $V(\mathbf{r})=2\pi/\kappa\delta(\mathbf{r})(1+\partial_{r})$ is the regularized atomic zero-range potential, $\mathbf{E}(t)$ the laser electric field, $\hat{\mathbf{E}}(\mathbf{r},t)=\sum_{\mathbf{k}}\sqrt{\frac{2\pi\omega}{\cal{V}}}\hat{\mathbf{e}}\left(a_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}-i\omega t}+a^{+}_{\mathbf{k}}e^{i\mathbf{k}\cdot\mathbf{r}-i\omega t}\right)$ is the field of the emitted photons in second quantization, with the creation and annihilation operators $a^{+}_{\mathbf{k}}$ and $a_{\mathbf{k}}$, the emitted photon frequency $\omega=c|\mathbf{k}|$, the polarization unit vector $\hat{\mathbf{e}}$, the speed of light $c$, and the quantization volume $\cal{V}$. Atomic units are used throughout. In the nonrelativistic regime the dipole approximation for the emission is applied, neglecting the spatial dependence of the emission field $\mathbf{k}\cdot\mathbf{r}\ll 1$.
With the electron wave function $\Psi$ as a solution of the Schrödinger equation in the atomic potential and the laser electric field, the amplitude of the coherent photon emission process in the first-order perturbation will read
$m_{\mathbf{k}}=-i\int ds\langle\Psi(s),1_{\mathbf{k}}|\mathbf{r}\cdot\hat{\mathbf{E}}(s)|\Psi(s),0_{\mathbf{k}}\rangle,$ (2)
yielding the differential probability for the coherent photon emission:
$\frac{dw}{d^{3}\mathbf{k}}=\left|m_{\mathbf{k}}\right|^{2}\frac{\cal{V}}{(2\pi)^{3}}.$ (3)
Describing the interaction of the electron with the photon field in the velocity gauge:
$m_{\mathbf{k}}=-i\int ds\langle\Psi(s),1_{\mathbf{k}}|\hat{\mathbf{A}}(s)\cdot\hat{\mathbf{p}}|\Psi(s),0_{\mathbf{k}}\rangle,$ (4)
where $\hat{\mathbf{A}}=-\hat{\mathbf{E}}(t)/\omega$ is the vector potential operator for the emitted photon.
After projecting on the photon initial and final states, and integrating over the emission direction, we obtain
$\frac{dw}{d\omega}=\frac{2\alpha\omega}{3\pi c^{2}}\left|\int ds\langle\Psi(s)\left|\hat{\mathbf{p}}\right|\Psi(s)\rangle\exp(i\omega s)\right|^{2}$ (5)
with the fine-structure constant $\alpha$. The spectral intensity of Eq. (5) can be viewed as classical radiation of the quantum current $\mathbf{j}(\mathbf{r},t)=\Psi(\mathbf{r},t)^{*}\hat{\mathbf{p}}\Psi(\mathbf{r},t)$. In this way the coherent part of the radiation is derived [24].
Figure 2: Coherent radiation spectra during tunneling ionization in the asymmetric laser pulse: (a,c) $E_{0}=0.1$ [zoom in Fig. 1(a)]; (b,d) $E_{0}=0.075$. Comparison of the Brunel radiation via $j_{11}$ with the three-step HHG ($j_{01}+j_{10}$) and with the Babushkin model ($j_{B}$). (c,d) Zoom in the low energy spectral region $\omega<0.2$ eV.
Figure 3: The role of the under-the-barrier dynamics for coherent radiation in a half-cycle pulse at $\omega=0.1\omega_{0}$ (1D model): The function $j_{\omega}(x)j^{*}_{\omega}(y)$ is plotted, which determines the coherent photon emission probability via Eq. (16). Blue line indicates the tunnel exit $x_{e}=\kappa^{2}/(2E_{0})$. The region inside the blue sector shows nonnegligible contribution to the radiation.
Figure 4: Spontaneous radiation spectrum during tunneling ionization: (a) asymmetric few-cycle laser pulse, (b) symmetric laser pulse, in both case $N=2$.
We describe the electron state $\Psi$ in Eq. (5) during tunneling ionization via the strong-field approximation (SFA) [25]:
$\Psi=\psi^{(0)}+\psi^{(1)}+\psi^{(2)},$ (6)
where the zeroth-order term $\psi^{(0)}$ coincides with the unperturbed bound state $\psi^{(a)}$ in the atomic potential. In the first-order SFA
$\psi^{(1)}=i\int^{s}dt\int d\mathbf{p}\psi^{f}_{\mathbf{p}}\langle\psi^{f}_{\mathbf{p}}(t)|\mathbf{r}\cdot\mathbf{E}(t)|\psi^{a}(t)\rangle$ (7)
with the unperturbed continuum state $\psi^{f}$ in the laser field, Volkov state [26], and in the second-order SFA [25]:
$\psi^{(2)}$ $=$ $i\int^{s}dr\int^{r}dt\int d\mathbf{p}\int d\mathbf{q}\psi^{f}_{\mathbf{p}}$ (8)
$\times\langle\psi^{f}_{\mathbf{p}}(r)|V|\psi^{f}_{\mathbf{q}}(r)\rangle\langle\psi^{f}_{\mathbf{q}}(t)|\mathbf{r}\cdot\mathbf{E}(t)|\psi^{a}(t)\rangle.$
The wave function $\psi^{(1)}$ describes the electron dynamics in the continuum, including the under-the-barrier dynamics, and the motion in the laser field after leaving the barrier. In Eq. (5) we can identify different probability currents $j(s)=j_{0,1}(s)+j_{1,0}(s)+j_{1,1}(s)$, where
$j_{0,1}(s)$ $=$ $\langle\psi^{(0)}(s)\left|\hat{p}_{x}\right|\psi^{(1)}(s)\rangle,$
$j_{1,0}(s)$ $=$ $\langle\psi^{(1)}(s)\left|\hat{p}_{x}\right|\psi^{(0)}(s)\rangle,$ (9)
$j_{1,1}(s)$ $=$ $\langle\psi^{(1)}(s)\left|\hat{p}_{x}\right|\psi^{(1)}(s)\rangle.$
The second-order terms $j_{0,2}$, $j_{2,0}$ describe the contribution to the radiation due to the under-the-barrier recollisions, which are usually small in the net radiation, see Sec. III. The radiation via the combination of $j_{01}$ and $j_{10}$ emerges due to the recombination to the bound state and, therefore, describes three-step HHG, including below-threshold harmonics. The radiation via $j_{11}$ term includes the effect of the electron transition to the continuum and the continuum dynamics, which will be the main point of our study.
For the physical interpretation, we introduce an heuristic model similar to Babushkin [10], to simulate the radiation only due to the continuum dynamics excluding the under-the-barrier contribution. In this model we evaluate the ionized electron probability current in the continuum, consisting of the tunnel ionized electron density via Amosov-Delone-Krainov (ADK) theory [27, 28] multiplied by the laser driven velocity:
$\mathbf{j}_{B}(s)=\int^{s}dt[\mathbf{A}(s)-\mathbf{A}(t)]\,\frac{\kappa^{2}}{2}\frac{E_{0}}{E_{a}}\exp\left[-\frac{2E_{a}}{3|\mathbf{E}(t)|}\right],$ (10)
with the atomic field $E_{a}=\kappa^{3}$, $\kappa=\sqrt{2I_{p}}$, and the atomic ionization energy $I_{p}$, and calculate the corresponding spectrum via Eq. (5).
We have carried out the calculation of the coherent radiation spectra in the case of a symmetric and asymmetric few-cycle laser pulses defined by their displacement
${\bm{\alpha}}(t)=-\hat{\mathbf{e}}_{x}\frac{E_{0}}{\omega_{0}^{2}}\left[\cos(\omega_{0}t)+\mathscr{A}\sin(2\omega_{0}t)\right]\cos(\omega_{0}t/N)^{4},$ (11)
where $\alpha^{\prime\prime}(t)=A^{\prime}(t)=-E(t)$, $E_{0}$ and $\omega_{0}$ are the laser field amplitude and frequency, respectively, $N=6$, and $\cal A$} is the asymmetry parameter of the laser pulse, ${\cal A}=0$ for the symmetric monochromatic pulse, and ${\cal A}=0.1$ for the asymmetric two-color one. In the calculation of the different currents $j_{ij}$, the time integrations are performed numerically, while all other integrals are evaluated analytically in the applied 3D zero-range potential case, using the corresponding ionization matrix element
$\int d^{3}\mathbf{r}\psi_{\mathbf{p}}(\mathbf{r})^{*}x\psi^{a}(\mathbf{r})=-\frac{2i\sqrt{\kappa}p_{x}}{\pi\left(\kappa^{2}+p^{2}\right)^{2}},$ (12)
where $\psi_{\mathbf{p}}$ is a plane wave with momentum $\mathbf{p}$. Further, boundary terms in the temporal Fourier transformation of the currents are treated as follows
$\int_{-\infty}^{\infty}f(s)ds=\frac{f(t_{i})^{2}}{f^{\prime}(t_{i})}+\int^{t_{f}}_{t_{i}}dsf(s)-\frac{f(t_{f})^{2}}{f^{\prime}(t_{f})}.$ (13)
The spectra in a symmetric and asymmetric few-cycle laser pulses with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=6$ in the 3D zero-range potential are shown in Fig. 1. In both cases the three-step HHG dominates after third harmonic, $\omega>4$ eV, while the continuum contribution via $j_{11}$ term does at low frequencies $\omega<4$ eV, see zoom of spectra for the case of the asymmetric laser pulse in Fig. 2. The first and second harmonics are mostly due to Thomson scattering, and the SFA calculation via $j_{11}$ is well approximated by the Babushkin model in the interval $0.5<\omega<4$ eV.
In the asymmetric pulse [Fig. 2], the radiation has a spike in the near-zero-frequency domain dominated by the continuum contribution ($j_{11}$). The latter is similar to the Babushkin model. The spike at small frequencies has a simple explanation. Initially, the electron is at rest at the core, then it tunnels out during the current formation time $\tau$, and is accelerated in the continuum due to the laser field up to its final velocity $v$. The current associated with this motion can be modeled as
$j(t)\sim v\,\,[1+\text{erf}(t/\tau)],$ (14)
with the Fourier component $j_{\omega}\sim v\exp(-\omega^{2}\tau^{2}/4)/\omega$. Thus, from Eq. (5)
$\frac{dw}{d\omega}\sim\frac{v^{2}}{\omega}e^{-\omega^{2}\tau^{2}/2},$ (15)
and the spectrum has a singularity at zero-frequency.
While $j_{11}$ shows the near-zero-frequency spike in the spectrum, it however includes additionally the under-the-barrier effect, which enhances the emission in most of the spectral region (at $\gamma=0.5$ the Babushkin model exaggerates the singularity because it does not account accurately the formation time of the current. Here, the error-function is replaced by a Heaviside-theta function). The enhancement is mostly due to the acceleration of the electron during tunneling. At larger Keldysh parameters $\gamma=\kappa\omega_{0}/E_{0}$ [cf. the cases $\gamma=0.5$ and $\gamma=0.7$ in Fig. 2] the effect of the under-the barrier dynamics is stronger. In particular, the enhancement of the near-zero frequency peak with respect to the Babushkin model is larger. However, the probability of the radiation is decreased. The width of the near-zero-frequency spike is about $\Delta\omega\sim 0.2$ eV, which is related to the laser pulse duration. In fact, the FWHM of the pulse in intensity is approximately $\tau\sim 200$ a.u. which corresponds to an energy of $1/\tau\sim 0.15$ eV, roughly fitting to $\Delta\omega\sim 1/\tau$.
To demonstrate the contribution of the under-the-barrier dynamics for the coherent radiation, we present the emission probability of Eq. (5) in the following form in the 1D case:
$\frac{dw}{d\omega}=\frac{2\alpha\omega}{3\pi c^{2}}\int dx\int dyj_{\omega}(x)j^{*}_{\omega}(y),$ (16)
with the current Fourier components $j_{\omega}(x)$. The function $j_{\omega}(x)j^{*}_{\omega}(y)$ is plotted in Fig. 3. From the latter we can distinguish which coordinate regions $x\sim y$ contribute significantly in the emission probability. We see in Fig. 3 that the under-the-barrier region has nonnegligible contribution to the radiation.
III Spontaneous radiation during tunneling ionization
The spontaneous radiation during tunneling ionization is calculated as a photon emission during the transition between the state of the tunneling electron $\Psi(t)$ and the continuum state in the laser field:
$dw=\left|m(\mathbf{k},\mathbf{p})\right|^{2}\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\frac{d^{3}\mathbf{p}}{(2\pi)^{3}},$ (17)
where the emission amplitude for the photon momentum $\mathbf{k}$, and the electron final momentum $\mathbf{p}$ reads
$m(\mathbf{k},\mathbf{p})=\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle,$ (18)
with the SFA wave function of the tunneling electron $\Psi(t)$. Here ${\cal V}=1$ is assumed. Integrating over the final electron momenta we obtain:
$\frac{dw}{d\omega}$ $=$ $\frac{4\pi\omega^{2}}{c^{3}}\int d^{3}\mathbf{p}\left|\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\Psi(t),0\rangle\right|^{2}.$ (19)
In the first-order SFA $\Psi=\psi^{(0)}(t)$ and we obtain the photon emission probability which coincides with the result of Ref. [14] in velocity gauge. In the second-order SFA calculation, the electron initial state during the photon emission process is $\Psi(t)=\psi^{(0)}(t)+\psi^{(1)}(t)$, which includes the continuum dynamics in the tunneling ionization process via $\psi^{(1)}(t)$. Comparing the first-order probability with the second-order one, we learn about the role of the continuum dynamics for the spontaneous radiation. We define the zero-order amplitude as
$m_{0}(\mathbf{k},\mathbf{p})=\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\psi^{(0)}(t),0\rangle,$ (20)
and the first-order one
$m_{1}(\mathbf{k},\mathbf{p})=\int dt\langle\psi_{\mathbf{p}}^{f}(t),1|\hat{\mathbf{p}}\cdot\hat{\mathbf{A}}(t)|\psi^{(1)}(t),0\rangle,$ (21)
as well as the total amplitude
$m(\mathbf{k},\mathbf{p})=m_{0}(\mathbf{k},\mathbf{p})+m_{1}(\mathbf{k},\mathbf{p}).$ (22)
We have carried out the calculation of the incoherent radiation spectra in the case of a symmetric and asymmetric few-cycle laser pulses defined above with $E_{0}=0.1$, $\omega_{0}=0.05$, and $N=2$.
Here again, in the calculation of the amplitudes $m_{i}$ the time integrations are performed numerically. All other inner integrals are calculated analytically. Especially, $m_{1}$ has the structure of a nested integral and is transformed via two partial integrations in the following way
$m_{1}$ $=$ $\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\exp(i\omega s)\int^{s}dt\tilde{m}_{1}(t)$ (23)
$=$ $-\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\tilde{m}_{1}(s)\int^{s}dt\exp(i\omega t)$
$+\int ds\mathbf{E}(s)\int^{s}dt\tilde{m}_{1}(t)\int^{s}dt\exp(i\omega t)$
$=$ $-\int ds\left[\mathbf{p}+\mathbf{A}(s)\right]\tilde{m}_{1}(s)\int^{s}dt\exp(i\omega t)$
$-\int ds\tilde{m}_{1}(s)\int^{s}dr\mathbf{E}(r)\int^{r}dt\exp(i\omega t).$
All inner time integrals are analytical and the outer numerical integration one-dimensional over the full time-space. The final momentum integration is also performed numerically. Further, boundary terms of the time integrals are treated like
$\int dsf(s)=\int^{t_{f}+2i\gamma/\omega}_{t_{i}+2i\gamma/\omega}dsf(s).$ (24)
The spectra for the spontaneous radiation are shown in Fig. 4. The spectrum via $m_{0}(\mathbf{k},\mathbf{p})$ is similar to the calculation of Ref. [14], with the spectral width $\Delta\omega\sim 1/\tau_{K}$, determined by the Keldysh time $\tau_{K}=\kappa/E_{0}$. Up to $\omega<6$ eV it is much smaller with respect to $m_{1}(\mathbf{k},\mathbf{p})$, which includes the continuum motion. In the near-zero-frequency region, $m_{1}(\mathbf{k},\mathbf{p})$ has a spike similar to that in the coherent radiation. At larger frequencies $\omega>6$ eV, $m_{0}(\mathbf{k},\mathbf{p})$ and $m_{1}(\mathbf{k},\mathbf{p})$ compensate each other. Thus, the amplitude $m_{0}(\mathbf{k},\mathbf{p})$ does not describe the observable spectrum neither at low frequencies $\omega\rightarrow 0$, nor at high ones $\omega\rightarrow\infty$, but only near the cutoff $\omega\sim 8$ eV. Our remarkable observation is that there is no significant difference in spontaneous radiation in symmetric or asymmetric pulses.
The spontaneous emission spectra in Fig. 4 have been calculated integrating over the final electron momentum. However, when the correlated electron and photon measurement is carried out in an experiment, one can access the electron momentum resolved photon spectra. In Fig.5 we present the electron momentum resolved total photon yield in the case of the correlated measurement. Calculation are carried out in the case of 1D potential and a half-cycle laser pulse, using saddle-point-approximation in evaluating the emission matrix element. Small contribution of the under-the-barrier recollisions via $m_{2}$ can be distinguished near the peak. However, this does not lead to any qualitative change of the spectrum.
Figure 5: Total probability of photon emission vs electron final momenta in the case of the correlated electron and photon measurement, for $E_{0}=0.2$ a.u. and a half-cycle pulse; the value for the $m_{0}$ contribution is multiplied by 2020.
IV Estimation of radiation intensities
Let us estimate the relation between the coherent and incoherent Brunel radiation. We use the dipole radiation power [29]:
$I=\frac{2\ddot{d}^{2}}{3c^{3}}.$ (25)
Estimating the current from Eq. (10):
$\dot{d}_{\rm coh}\sim(E_{0}/\omega_{0})w_{i}T,$ (26)
where $w_{i}\sim(\kappa^{2}/2)(E_{a}/E_{0})\exp[-2E_{a}/(3E_{0})]$ is the laser ionization rate, $T=2\pi/\omega_{0}$ is the laser period. The probability of the coherent Brunel photon emission per laser period with the energy $\omega$ is
$w_{\rm coh}\sim\frac{I_{\rm coh}T}{\omega}\sim\frac{\alpha E_{0}^{2}}{c^{2}\omega}\left(\frac{\omega}{\omega_{0}}\right)^{2}w_{i}^{2}T^{3}\sim\alpha a_{0}^{2}\gamma^{-1}(w_{i}T)^{2},$ (27)
where $a_{0}=E_{0}/c\omega_{0}$ is the relativistic invariant field parameter.
Now let us estimate the probability of the spontaneous Brunel radiation again via the dipole radiation theory. For spontaneous radiation, we estimate the dipole moment derivative as
$\dot{d}_{\rm sp}\sim\omega_{c}\ell_{c},$ (28)
with the characteristic frequency $\omega_{c}\sim 1/\tau_{K}$, and $\ell_{c}\sim\kappa\tau_{K}$, with the Keldysh time $\tau_{K}=\kappa/E_{0}$, and the atomic velocity $\kappa$. Then, the spontaneous radiation power is
$I_{\rm sp}\sim\frac{\alpha\omega_{c}^{4}\ell^{2}}{c^{2}}w_{i}T,$ (29)
where we multiplied by the ionization probability because the dipole moment $d_{\rm sp}$ is created during tunneling. Then, the probability of the spontaneous photon emission during tunneling ionization:
$w_{\rm sp}\sim\frac{I_{\rm sp}T}{\omega_{c}}w_{i}T\sim\frac{\alpha\omega_{c}^{3}\ell^{2}}{c^{2}}w_{i}T^{2}\sim\alpha\frac{\kappa^{2}}{c^{2}}\gamma^{-1}w_{i}T,$ (30)
with the Keldysh parameter $\gamma=\omega_{0}\tau_{K}$.
Thus, the ratio of the probabilities is
$\frac{w_{\rm sp}}{w_{\rm coh}}\sim\frac{\kappa^{2}}{c^{2}}\frac{1}{a_{0}^{2}w_{i}T}$ (31)
For $E_{0}=0.1$ (the laser intensity $3.5\times 10^{14}$ W/cm2), $E_{0}/E_{a}\sim 0.1$, and $a_{0}\sim 1.5\times 10^{-2}$, $\frac{I_{p}}{\omega_{0}}\sim 10$, $e^{-\frac{2E_{a}}{3E_{0}}}\sim 1.3\times 10^{-3}$, $w_{i}T\sim 10^{-2}$, we obtain $w_{\rm sp}/w_{\rm coh}\sim 10^{2}$, which is roughly in accordance with the spectra of Figs. 1 and 4.
We may estimate also the spontaneous radiation per electron: $w_{\rm sp}^{(e)}\sim w_{\rm sp}/w_{i}T\sim(\alpha/\gamma)(\kappa/c)^{2}\sim 10^{-6}$, which indicates complications for the electron and photon coincidence measurement. The Brunel radiation scales as $1/\gamma$ with the Keldysh parameter.
When the Brunel radiation is observed with a gas target, the coherent Brunel radiation will be dominating over the spontaneous one, if the target density is large enough: ${\cal N}=n{\cal V}>w_{\rm sp}/w_{\rm coh}$, with the particle number density $n$, the focal volume ${\cal V}\sim(\pi w^{2}/\lambda)(\pi w^{2})$, and the laser beam waist size $w$. With $w\sim 3\lambda$, $\lambda\sim 1\,\mu$, this will take place at a gas density $n>10^{11}$ cm-3.
V Conclusion
We have investigated the coherent and incoherent radiation during tunneling ionization. The coherent radiation in the near-zero-frequency ($\omega\sim 0.2$ eV) domain is dominated by the continuum dynamics via the $j_{11}$ term in the current. We have shown that due to the under-the-barrier motion the radiation can be enhanced. The nonadiabaticity increases the effect of the under-the-barrier dynamics at large Keldysh parameters, however generally decreasing the radiation yield.
The spontaneous radiation is not affected by the asymmetry of the driving laser pulse. The characteristic width of the spontaneous radiation is determined by the Keldysh time $\Delta\omega\sim 1/\tau_{K}$. There are two contributions into the spontaneous radiation: due to the direct transition during tunneling (with the amplitude $m_{0}$) investigated in Ref [14], and due to additional motion in the continuum (with the amplitude $m_{1}$). At larger frequencies $\omega\gtrsim 1/\tau_{K}$, $m_{0}$ and $m_{1}$ terms compensate each other and the total radiation is suppressed. At low frequencies $\omega\lesssim 1/\tau_{K}$, the radiation is dominated by the continuum dynamics via the amplitude $m_{1}$. Thus, the contribution of the direct transition in the radiation, corresponding to the result of Ref. [14], essentially describes the spontaneous radiation only near the cutoff.